### v1

# Non-Gaussianities in primordial black hole formation and induced gravitational waves

Abstract
The most promising mechanism of generating primordial black holes (PBHs) is by the enhancement of power spectrum of the primordial curvature perturbation, which is usually accompanied by the the enhancement of non-Gaussianity that crucially changes the abundance of PBHs. In this review I will discuss how non-Gaussianity is generated in single field inflation as well as in the curvaton scenario, and then discuss how to calculate PBH mass function with such non-Gaussianities. I also show non-Gaussianity only has mild effects on the induced gravitational waves (GWs), which gives robust predictions in the mHz and nHz GW experiments.
I Overview
[…]
A related phenomenon is the stochastic gravitational wave background induced by the enhanced curvature perturbation. In General Relativity, the curvature perturbation $\mathcal{R}$ and tensor perturbation $h_{ij}$ are only coupled at nonlinear order, and the most important coupling is the tensor-scalar-scalar type $\sim h_{ij}\partial_{i}\mathcal{R}\partial_{j}\mathcal{R}$ Matarrese et al. (1993, 1994, 1998); Noh and Hwang (2004); Carbone and Matarrese (2005); Nakamura (2007); Ananda et al. (2007); Osano et al. (2007); Baumann et al. (2007). Therefore, the curvature perturbation can source the GWs after it reenters the Hubble horizon, which is called scalar induced secondary GWs. The peak frequency, as well as the PBH mass generated by such a perturbation, are determined by the horizon scale at the reentry moment, which are connected by Saito and Yokoyama (2009)
$f_{\mathrm{IGW}}\approx 3\mathrm{Hz}\left(\frac{M_{\mathrm{PBH}}}{10^{16}\mathrm{g}}\right)^{-\frac{1}{2}}.$ (1)
At the peak scale, the spectral energy density of the induced GWs we observe today (normalized by the critical density), is roughly $\Omega_{\text{GW,0}}\equiv\rho_{\mathrm{cr}}^{-1}\left(\mathrm{d}\rho_{\mathrm{GW}}/\mathrm{d}\ln k\right)_{t_{0}}\sim 10^{-6}\mathcal{P}_{\mathcal{R}}^{2}$, where $\mathcal{P}_{\mathcal{R}}$ is the primordial power spectrum of $\mathcal{R}$. The nearly scale-invariant power spectrum extrapolated from CMB scales predicts a scale-invariant induced GW with an amplitude of $\Omega_{\text{GW,0}}\sim 10^{-24}$, far less than observational bound Baumann et al. (2007). However, for the enhanced curvature perturbation with abundant PBH formation ($\mathcal{P_{R}}\sim\mathcal{O}(10^{-2})$), the induced GW is also enhanced to $\Omega_{\text{GW,0}}\sim 10^{-10}$, which reaches the detectable range of many experiments like the next-generation ground- and space-borne interferometers and pulsar timing arrays. As both $\Omega_{\text{GW,0}}$ and PBH abundance $f_{\mathrm{PBH}}$ depend on $\mathcal{P_{R}}$, they can be cross-checked to probe the primordial power spectrum on small scales Saito and Yokoyama (2009, 2010); Bugaev and Klimai (2010); Assadullahi and Wands (2010); Bugaev and Klimai (2011). For more discussion on the induced GWs, see Domènech (2024).
However, to enhance the power spectrum of the curvature perturbation, usually the slow-roll attractor is deviated, while the shape and amplitude of the enhanced power spectrum, as well as the statistics are also changed, depending on the model. Even a small non-Gaussianity can change the relation between the PBH abundance and induced GW spectrum significantly, as the induced GW depends mainly on the variance of the PDF of the curvature perturbation, while the PBH abundance depends on the high-$\sigma$ tail of the PDF. A small change in the PDF, i.e. deviation of the Gaussianity, can give rise to a huge enhancement/suppression in the PBH abundance Bullock and Primack (1997); Ivanov (1998); Yokoyama (1998); Pina Avelino (2005); Seery and Hidalgo (2006); Hidalgo (2007); Byrnes et al. (2012); Young and Byrnes (2013); Young et al. (2016); Atal and Germani (2019); Yoo et al. (2019); Kehagias et al. (2019); Mahbub (2020); Riccardi et al. (2021); Taoso and Urbano (2021); Biagetti et al. (2021); Kitajima et al. (2021); Young (2022); Escrivà et al. (2022a); Matsubara and Sasaki (2022); Gow et al. (2022), but only change the amplitude of the induced GW spectrum mildly Nakama et al. (2017); Garcia-Bellido et al. (2017); Cai et al. (2019a); Unal (2019); Unal et al. (2020); Adshead et al. (2021); Garcia-Saenz et al. (2022); Abe et al. (2022).
On large scales, non-Gaussianity is often described by the perturbative series Komatsu and Spergel (2001)
$\mathcal{R}=\mathcal{R}_{g}+\frac{3}{5}f_{\mathrm{NL}}(\mathcal{R}_{g}^{2}-\langle\mathcal{R}_{g}\rangle^{2})+\cdots,$ (2)
which was used to describe the curvaure perturbation on small scales for PBH formation Byrnes et al. (2012); Young and Byrnes (2013); Tada and Yokoyama (2015); Young and Byrnes (2015); Young et al. (2016); Franciolini et al. (2018); Ando et al. (2018a); Atal and Germani (2019); Passaglia et al. (2019) as well as the induced GWs Nakama et al. (2017); Garcia-Bellido et al. (2017); Cai et al. (2019a); Unal (2019); Unal et al. (2020). However, in many models, the non-linear parameter $f_{\mathrm{NL}}$ reaches $\mathcal{O}(1)$ Namjoo et al. (2013); Martin et al. (2013); Chen et al. (2013); Motohashi et al. (2015); Davies et al. (2022); Namjoo (2023); Namjoo and Nikbakht (2024), so the higher orders in (2) might be important, which are highly model-dependent. These higher order terms must be taken into account in the PBH formation. Fortunately, by $\delta N$ formalism, a fully nonlinear curvature perturbation can be derived for single field inflation and for the curvaton scenario, which indeed alters the PDF of $\mathcal{R}$ completely in large $\mathcal{R}$ regime. Various methods are developed to calculate the PBH abundance for such a PDF, including the Press-Schechter-type formalism Biagetti et al. (2021); Gow et al. (2022); Ferrante et al. (2022) and the peak theory Green et al. (2004); Yoo et al. (2018); Germani and Musco (2019); Atal et al. (2019, 2020a); Young and Musso (2020); Yoo et al. (2021); Taoso and Urbano (2021); Riccardi et al. (2021); Kitajima et al. (2021); Young (2022). The Press-Schechter-type formalism for Gaussian curvature perturbation is straightforward to be further extended to non-Gaussian case.
Considering non-Gaussianity on the PBH formation and induced GW generation is important in many theoretical issues and observational problems in cosmology.
  • PBHs might be the supermassive or stupendously large BHs which seed the galaxy or structure formation Bean and Magueijo (2002); Kawasaki et al. (2012); Carr and Silk (2018); Carr et al. (2021). To satisfy the stringent constraints from CMB $\mu$- and $y$-distortion, large non-Gaussianities must be included Nakama et al. (2016, 2018, 2019); Atal et al. (2020b); Carr and Silk (2018); Liu and Bromm (2022); Biagetti et al. (2023); Gouttenoire et al. (2023); Hooper et al. (2023).
  • According to the current observational constraints, asteroid-mass PBH can be all the dark matter, of which the concomitant induced GWs are detectable by space-borne interferometers LISA Barausse et al. (2020); Arun et al. (2022); Auclair et al. (2023); Bagui et al. (2023), Taiji Ren et al. (2023), and TianQin Liang et al. (2021). As positive non-Gaussianity can enhance the PBH abundance greatly, the required power spectrum of the curvature perturbation for $\Omega_{\mathrm{PBH}}=\Omega_{\mathrm{DM}}$ is reduced with positive non-Gaussianity, and the induced GW spectrum is also reduced, yet it is still above the LISA/Taiji/TianQin sensitivity curve Cai et al. (2019a); Bartolo et al. (2019); Unal (2019). The large non-Gaussian limit with $f_{\mathrm{NL}}\gg 1$ (i.e. the $\chi^{2}$-distribution) can be realized in the curvaton model when curvaton is negligible at its decay Bartolo et al. (2004); Enqvist and Nurmi (2005); Sasaki et al. (2006); Pi and Sasaki (2023a); Ferrante et al. (2022).
  • In 2023, the Pulsar Timing Array collaborations NANOGrav Agazie et al. (2023a, b), EPTA combined with InPTA Antoniadis et al. (2023a, b, c), PPTA Zic et al. (2023); Reardon et al. (2023a, b), CPTA Xu et al. (2023), as well as IPTA Agazie et al. (2023c), announced the strong evidence of nHz stochastic GW background, while the induced GW interpretation is favored Afzal et al. (2023). However, as the amplitude of the nHz GW spectrum is relatively high ($\sim 10^{-9}$), sub-solar-mass PBHs might be overproduced, i.e. $\Omega_{\mathrm{PBH}}>\Omega_{\mathrm{DM}}$ This tension can be solved by introducing a negative non-Gaussianity with $f_{\mathrm{NL}}\lesssim-1$ to suppress the PBH formation Franciolini et al. (2023a); Liu et al. (2023). Such a large and negative $f_{\mathrm{NL}}$ should be seen as the leading order of a fully nonlinear curvature perturbation, which can be realized in the curvaton scenario when the curvaton dominates at its decay moment Pi and Sasaki (2023a); Ferrante et al. (2022); Chen et al. (2023).
  • It was shown that PBHs obey a Poisson distribution when the density perturbation is Gaussian Ali-Haïmoud (2018); Desjacques and Riotto (2018), while non-Gaussianity can make PBHs to cluster initially on small scales Tada and Yokoyama (2015); Young and Byrnes (2015); Franciolini et al. (2018); Young and Byrnes (2020); Suyama and Yokoyama (2019); De Luca et al. (2021), which enhances the merger rate of PBHs significantly Clesse and García-Bellido (2017); Raidal et al. (2017). Such clustered PBHs can escape from the microlensing detections, which assume Poisson distribution García-Bellido and Clesse (2018).
  • When the power spectrum of the curvature perturbation is enhanced on small scales, the loop corrections on CMB scales might be large, which may jeopardize the mechanism of enhancing the power spectrum on small scales to generate PBHs Kristiano and Yokoyama (2022, 2023). It was soon realized that the loop correction can be suppressed when the transition to slow-roll inflation is smooth Riotto (2023); Firouzjahi (2023); Firouzjahi and Riotto (2024); Firouzjahi (2024a), as it is proportional to $f_{\mathrm{NL}}$. For recent discussions, see Motohashi and Tada (2023); Fumagalli (2023); Cheng et al. (2023); Franciolini et al. (2023b); Fumagalli et al. (2023); Maity et al. (2023); Iacconi et al. (2023); Firouzjahi (2024b).
For more discussion on the observational implications, see Part V and the references therein, for instance the review papers Carr et al. (2020); Carr and Kuhnel (2020); Carr et al. (2024).
This paper is organized as follows. In Section II we review calculation of PBH mass function, emphasizing on when and how the non-Gaussianity comes into play. In Section III, we study some typical inflation models which can generate large non-Gaussianity: the ultra-slow-roll inflation, constant-roll inflation with a bumpy potential, inflation with piecewise quadratic potential, and the curvaton scenario. After getting the non-Gaussian curvature perturbations in these models, we calculate the PBH mass function and abundance by the Press-Schechter-type formalism in Section IV, as well as the scalar-induced GWs in Section V. We conclude by discussing some interesting topics to be studied in the future in Section VI.
II A brief review of PBH formation
II.1 PBH abundance in the Gaussian case
In this section we will briefly review how to calculate the PBH mass function and abundance by the Press-Schechter formalism, emphasizing on where and how the non-linearities and primordial non-Gaussianities enter. Our starting point is the power spectrum of the primordial curvature perturbation on comoving slices, $\mathcal{P_{R}}(k)$, generated at the end of inflation. The simplest Press-Schechter formalism uses the density contrast on comoving slices to calculate, which is related to $\mathcal{R}$ by the Poisson equation
$\delta\equiv\left(\frac{\delta\rho}{\rho_{0}}\right)_{c}=-\frac{2(1+w)}{5+3w}\frac{1}{H^{2}a^{2}}\nabla^{2}\mathcal{R}.$ (3)
In radiation dominated case ($w=1/3$), the prefactor is 4/9.
The density contrast threshold $\delta_{\mathrm{th}}$ is to determine whether a PBH could form in a specific Hubble patch or not, which was argued by the Jeans instability to be approximately $\delta_{\text{th}}\approx w$ Carr (1975). A recent analytical result is the Harada-Yoo-Kohri (HYK) threshold Harada et al. (2013):
$\delta_{\text{th}}=\frac{3(1+w)}{5+3w}\sin^{2}\left(\frac{\pi\sqrt{w}}{1+3w}\right)\xrightarrow{\quad w\to 1/3\quad}0.41.$ (4)
Given the PDF of the curvature perturbation $\mathbb{P}(\mathcal{R})$, the PDF of the density contrast is given by
$\mathbb{P}(\delta)=\mathbb{P}(\mathcal{R})\left|\frac{\mathrm{d}\mathcal{R}}{\mathrm{d}\delta}\right|,$ (5)
where the Jacobian $|\mathrm{d}\mathcal{R}/\mathrm{d}\delta|$ can be easily calculated from the linear Poisson equation (3). With these quantities, we can calculate the PBH abundance in the following way:
[…]
We will not discuss the issues of window function in this paper. The key point of the above calculation is that, if the curvature perturbation $\mathcal{R}$ is Gaussian, the density contrast is also Gaussian, guaranteed by the linear relation (3). Therefore the integration (6) is simply a complementary error function, which gives a Gaussian suppression $\sim\exp(-\delta^{2}/2\delta_{\mathrm{th}}^{2})$ at the large tail.
II.2 Press-Schechter-type formalism
The estimation (6) neglect some important facts in the calculation of PBH abundance, which might change the result significantly.
II.2.1 Compaction function
The criterion of PBH formation is not simply a threshold of the density contrast $\delta\rho/\rho$ Numerical study on the black hole collapse show that the criterion should be put on the averaged mass excess inside a sphere of a fixed areal radius, dubbed the compaction function Shibata and Sasaki (1999). In a reduced-circumference polar coordinates, we have  [^2] In this paper, following Escrivà et al. (2022b); Germani and Sheth (2023), we write the spatial metric as $\mathrm{d}\Sigma^{2}=e^{2\mathcal{R}(r)}\left(\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega^{2}\right)=\frac{\mathrm{d}\bar{r}^{2}}{1-K(r)\bar{r}^{2}}+\bar{r}^{2}\mathrm{d}\Omega^{2}$ (7) where $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}$ is the two-dimensional line element on a sphere. $\bar{r}$ is called reduced-circumference radius, which is related to $r$ by $\bar{r}=e^{\mathcal{R}(r)}r.$ (8)
$\mathscr{C}(t,\bar{r},\mathbf{x})\equiv\frac{2\delta M(r,t,\mathbf{x})}{r}=\frac{3}{\bar{r}}\left(aH\right)^{2}\int^{\bar{r}}_{0}\mathrm{d\varrho}~{}\varrho^{2}\delta(t,\varrho,\mathbf{x}).$ (9)
We can see that the compaction function is the averaged density contrast inside a sphere of radius $\bar{r}$, which is conserved on superhorizon scales. For the discussion of gauge issue, see Harada et al. (2015). It seems that, if we substitute the linear relation (3) into (9), we have
$\mathscr{C}_{\ell}=-\frac{6(1+w)}{5+3w}r\frac{\partial\mathcal{R}}{\partial r}\xrightarrow{\quad w=1/3\quad}-\frac{4}{3}r\frac{\partial\mathcal{R}}{\partial r}.$ (10)
However, linear relation (3) is only the leading-order approximation of Harada et al. (2015); Kawasaki and Nakatsuka (2019); Young et al. (2019); De Luca et al. (2019)
$\delta=-\frac{4}{3}\cdot\frac{3(1+w)}{5+3w}\left(\frac{1}{aH}\right)^{2}e^{-5\mathcal{R}/2}\nabla^{2}e^{\mathcal{R}/2},$ (11)
Therefore we should use
$\mathscr{C}=\mathscr{C}_{\ell}\left(1-\frac{5+3w}{12(1+w)}\mathscr{C}_{\ell}\right)\xrightarrow{\quad w=1/3\quad}\mathscr{C}=\mathscr{C}_{\ell}-\frac{3}{8}\mathscr{C}_{\ell}^{2}.$ (12)
Such a nonlinear relation will naturally bring non-Gaussianity to the calculation of PBH formation even if $\mathcal{R}$ is Gaussian. From now on, we will focus on radiation-dominated universe with $w=1/3$.
II.2.2 Threshold
Around a density peak, the compaction function $\mathscr{C}(r)$ is a function of the radius $r$. Numerical calculations tell us that whether an overdensed region can collapse into a black hole depends on whether the maximum of the compaction function, $\mathscr{C}(r_{m})$, exceeds the critical value $\mathscr{C}_{\mathrm{th}}$, which varies from 0.4 to 0.67, depending on the density profile of the overdensed region. It was further found that only the width of the compaction function peak, parameterized by $q\equiv-\mathscr{C}^{\prime\prime}(r_{m})r_{m}^{2}/(4\mathscr{C}(r_{m}))$, is important Musco (2019). Large $q$ corresponds to a narrow $\mathscr{C}(r_{m})$, given by a top-hat density profile which drops rapidly around $r_{m}$. Strong pressure gradients at $r_{m}$ can resist the gravitational collapse, thus $\mathscr{C}_{\mathrm{th}}$ reaches its maximum $2/3$ in this limit. On the contrary, small $q$ corresponds to a broad $\mathscr{C}(r_{m})$, given by a density profile highly peaked at the center. As the pressure gradients can be neglected, it degenerate to the HYK limit $\mathscr{C}_{\mathrm{th}}\approx\delta_{\mathrm{th}}\approx 0.4$. In principle the density profiles and their statistical properties should be input if we want to calculate PBH abundance appropriately. Some hypothetical profiles are given by fitting the numerical results Musco (2019); Escrivà et al. (2020); Young (2019). In the peak theory, the profiles can be derived by the multipole moments of the power spectrum, provided the curvature perturbation is Gaussian Yoo et al. (2018); Atal et al. (2019, 2020a); Yoo et al. (2021); Kitajima et al. (2021).
By using (12), $\mathscr{C}^{\prime}(r)=0$ gives $\mathscr{C}_{\ell}^{\prime}(r)=0$ or $\mathscr{C}_{\ell}(r)=4/3$. The former condition determines the extreme of $\mathscr{C}(r)$ at $r_{m}$, where its second derivative is
$\mathscr{C}^{\prime\prime}(r_{m})=\left(1-\frac{3}{4}\mathscr{C}_{\ell}(r_{m})\right)\mathscr{C}_{\ell}^{\prime\prime}(r_{m}).$ (13)
We see that depending on whether $\mathscr{C}_{\ell}<4/3$ or $\mathscr{C}_{\ell}>4/3$, the maximum $\mathscr{C}(r_{m})$ corresponds to the maximum or minimum of $\mathscr{C}_{\ell}$. Only the former case, dubbed type I PBH Kopp et al. (2011), is consistent with the linear approximation. Type II PBHs with $\mathscr{C}_{\ell}>4/3$ are rarer, which we do not consider in this paper. Given the profile-dependent threshold on $\mathscr{C}$, the corresponding threshold on the linear compaction function $\mathscr{C}_{\ell}$ can be solved by (12)
$\mathscr{C}_{\ell,\mathrm{th}}=\frac{4}{3}\left(1-\sqrt{1-\frac{3}{2}\mathscr{C}_{\mathrm{th}}}\right).$ (14)
As two examples, the threshold of a top-hat profile $\mathscr{C}_{\mathrm{th}}\approx 1/2$ gives $\mathscr{C}_{\ell,\mathrm{th}}\approx 2/3$. In the peak theory, a monochromatic power spectrum gives $\mathscr{C}_{\mathrm{th}}\approx 0.587$, thus $\mathscr{C}_{\ell,\mathrm{th}}\approx 0.872$ Kitajima et al. (2021).
When the power spectrum of $\mathcal{R}$ is broad, for every scale we study, we should add a window function to smooth out the unwanted short-wavelength fluctuations. This is highly non-trivial and there will be uncertainties from the choice of the window function, which will not be discussed in this paper.
II.2.3 Primordial non-Gaussianity
The primordial non-Gaussianity of $\mathcal{R}$ (if any) should be taken into account when deriving the PDF of $\mathscr{C}$ from the PDF of $\mathcal{R}$, which is the main topic of this paper.
We already know from Section II.2.1 that the compaction function $\mathscr{C}$ is a quadratic function of $\mathscr{C}_{\ell}$. In the presence of primordial non-Gaussianity in $\mathcal{R}$, even $\mathscr{C}_{\ell}$ is non-Gaussian. As we will see, in typical cases, $\mathcal{R}$ is a function of a Gaussian-distributed variable, $\mathcal{R}_{\mathrm{g}}$, which obeys the Gaussian PDF
$\mathbb{P}(\mathcal{R}_{\mathrm{g}})=\frac{1}{\sqrt{2\pi\sigma^{2}_{\mathcal{R}}}}\exp\left(-\frac{\mathcal{R}_{g}^{2}}{2\sigma_{\mathcal{R}}^{2}}\right),$ (15)
with
$\sigma_{\mathcal{R}}^{2}=\int\frac{\mathrm{d}k}{k}\tilde{W}^{2}(k,r)\mathcal{P}_{\mathcal{R}_{g}}(k).$ (16)
For simplicity, let’s assume that the non-Gaussianity is local-type, which gives
$\mathcal{R}=\mathscr{F}[\mathcal{R}_{\mathrm{g}}(r)],\qquad\mathcal{R}_{\mathrm{g}}=\mathscr{G}[\mathcal{R}(r)],$ (17)
where $\mathscr{G}\equiv\mathscr{F}^{-1}$ with $\mathscr{F}(0)=0$ and $\mathscr{F}^{\prime}(0)=1$, such that (17) can be expanded as $\mathscr{F}(\mathcal{R}_{\mathrm{g}})\approx\mathcal{R}_{\mathrm{g}}$ when $\mathcal{R}_{\mathrm{g}}$ is small. Concrete examples of such nonlinear relations (17) will be discussed in the next section. From (15), the corresponding PDF of $\mathcal{R}$ is
$\mathbb{P}(\mathcal{R})=\mathbb{P}_{\mathrm{G}}(\mathcal{R}_{\mathrm{g}})\cdot\left|\frac{\mathrm{d}\mathscr{F}}{\mathrm{d}\mathcal{R}_{\mathrm{g}}}\right|^{-1}=\mathbb{P}_{\mathrm{G}}(\mathcal{R}_{\mathrm{g}})\cdot\left|\frac{\mathrm{d}\mathscr{G}}{\mathrm{d}\mathcal{R}}\right|.$ (18)
This brings non-Gaussianity to the linear compaction function by
$\mathscr{C}_{\ell}=-\frac{4}{3}r\frac{\mathrm{d}\mathscr{F}}{\mathrm{d}\mathcal{R}_{\mathrm{g}}}\frac{\partial\mathcal{R}_{\mathrm{g}}(r)}{\partial r}.$ (19)
II.3 PBH abundance with primordial non-Gaussianity
According to the discussion above, the PBH abundance should be calculated as following Biagetti et al. (2021); Gow et al. (2022); Ferrante et al. (2022)
[…]
where PNG stands for primordial non-Gaussianity of the curvature perturbation generated during inflation, which will be discussed in details in the next section. Calculations following the routines in (20) are called extended Press-Schechter or Press-Schechter-type formalism, which is in principle different from another frequently used method, the peak theory Green et al. (2004); Yoo et al. (2018); Germani and Musco (2019); Atal et al. (2019, 2020a); Young and Musso (2020); Yoo et al. (2021); Taoso and Urbano (2021); Riccardi et al. (2021); Kitajima et al. (2021); Young (2022), by two main points. In extended Press-Schechter fomalism, the PBH abundance is calculated by the integrating the PDF of $\mathscr{C}_{\ell}$ from the threshold given by $\mathscr{C}_{\ell}\geq\mathscr{C}_{\ell,\mathrm{th}}$, while in the peak theory the PBH abundance is given by counting the number of peaks $n_{\mathrm{pk}}$ where the averaged compaction function in a sphere of radius $r_{m}$, $\langle\mathscr{C}\rangle_{r\leq r_{m}}$, exceeds a universal threshold 2/5 Escrivà et al. (2020).
III Primordial non-Gaussianities of the curvature perturbation
Usually, the primordial curvature perturbation on comoving slices $\mathcal{R}$ originates from the quantum fluctuations of the inflaton $\varphi$ during inflation Brout et al. (1978); Guth (1981); Starobinsky (1980); Mukhanov and Chibisov (1981); Linde (1982); Albrecht and Steinhardt (1982). In linear perturbation theory, $\mathcal{R}_{g}=-(H/\dot{\varphi})\delta\varphi$ Mukhanov (1985); Sasaki (1986), where $\delta\varphi$ is the field perturbation on spatially-flat slices. The non-Gaussian curvature perturbation is usually parameterized as the perturbative series of (17) Komatsu and Spergel (2001),
$\mathcal{R}=\mathcal{R}_{g}+\frac{3}{5}f_{\mathrm{NL}}\left(\mathcal{R}_{g}^{2}-\left\langle\mathcal{R}_{g}^{2}\right\rangle\right)+\cdots.$ (21)
In slow-roll inflation, the nonlinear parameter $f_{\mathrm{NL}}$ is proportional to the spectral tilt, i.e. the consistency relation $f_{\mathrm{NL}}=(5/12)(n_{s}-1)$ Maldacena (2003); Creminelli and Zaldarriaga (2004), which is suppressed by slow-roll parameters as $n_{s}-1=-2\epsilon_{H}-\eta_{H}$. On scales larger than 1 Mpc, the observational constraint is $f_{\mathrm{NL}}=-0.9\pm 5.1$ Akrami et al. (2020b); Abbott et al. (2022).
As is commented above, the PBH abundance depends sensitively on the non-Gaussianities, thus even a small non-Gaussianity can change the PBH abundance significantly. Furthermore, as the power spectrum is enhanced rapidly (usually as power-law) before the peak scale, an intuitive guess based on the consistency relation $f_{\mathrm{NL}}\sim n_{s}-1$ is that non-Gaussianity will also be enhanced to $\mathcal{O}(1)$. For instance, in ultra-slow-roll inflation, $\epsilon_{H}\lll 1$ and $\eta_{H}=-6$. If we only focus on the slow-roll attractor mode, we have
$f_{\mathrm{NL}}=\frac{5}{12}(n_{s}-1)=-\frac{5}{12}(2\epsilon_{H}+\eta_{H})\approx-\frac{5}{12}\times(-6)=\frac{5}{2}.$ (22)
Astonishingly, it gives the correct value, although it is in contradiction with the consistency relation of the total power spectrum $\mathcal{P_{R}}(k)\propto k^{4}$, of which the slope originates from the next-to-leading order of gradient expansion Leach et al. (2001); Byrnes et al. (2019); Cole et al. (2022).
In principle, when $f_{\mathrm{NL}}\gtrsim 1$, the higher order terms includes an infinite number of parameters, which makes it impossible to have a model-independent analysis  [^3] In some specific models, it is possible to truncate the perturbative series at quadratic level, no matter how large $f_{\mathrm{NL}}$ is. See Section III.4 for details.. Fortunately, the non-Gaussian curvature perturbation can be calculated by $\delta N$ formalism, based on the separate universe approach Sasaki and Stewart (1996); Wands et al. (2000); Lyth et al. (2005). During inflation, the gradient terms decay rapidly on superhorizon scales, which guarantees that in a region slightly larger than the Hubble patch, the inflaton field takes a homogeneous value, which varies from region to region. The initial condition for $\varphi$ in the numerous separate universes obeys a Gaussian distribution, which evolves independently on superhorizon scales. $\delta N$ formalism tells us that the final curvature perturbation in such a patch is the difference of total local expansion, i.e. the $e$-folding number $N$, between this patch to the fiducial $\langle N\rangle$, evaluated from an initial spatially-flat slice to the final comoving slice:
$\mathcal{R}=\delta N\equiv N(\varphi,\dot{\varphi})-\langle N(\langle\varphi\rangle,\langle\dot{\varphi}\rangle)\rangle.$ (23)
This formula does not require a decomposition of a background and a perturbation, thus can be used for large perturbations. We will use it to study the fully nonlinear curvature perturbation in the following sections. In Vennin and Wands (2024), (23) will be used to study the quantum diffusion.
III.1 Ultra-slow-roll inflation on a flat potential
When $f_{\mathrm{NL}}\gtrsim 1$, we should take into account all the higher order terms, which is only possible for some specific models. The simplest example is the so-called ultra-slow-roll inflation. Consider a plateau of potential $V(\varphi)=V_{0}$ for $\varphi_{i}<\varphi<\varphi_{t}$. On the plateau, the equation of motion of $\varphi$ is $\ddot{\varphi}+3H\dot{\varphi}=0$, where $H\approx\sqrt{V_{0}}/(\sqrt{3}M_{\mathrm{Pl}})$ is nearly a constant on the plateau. Conventionally, we define the $e$-folding number $N$, counted backward in time from $N(\varphi_{t})=0$, such that $\mathrm{d}N=-H\mathrm{d}t$. The equation of motion becomes
$\frac{\mathrm{d}^{2}\varphi}{\mathrm{d}N^{2}}-3\frac{\mathrm{d}\varphi}{\mathrm{d}N}=0.$ (24)
Two general solutions to (24) are $\varphi=\mathrm{constant}$ and $\varphi\propto e^{3N}$. The former static solution is the attractor, but it should not dominate at the classical level, otherwise the inflaton gets stuck on the plateau, which can only end with the help of quantum diffusion. The non-attractor decelerating solution $\varphi\propto e^{3N}$ is necessary as it helps the inflaton to classically shoot out of the potential plateau with large enough initial velocity. We will see that the excitation of such non-attractor solutions is the origin of the large non-Gaussianity. Together with the boundary condition at the endpoint $\varphi_{t}$ of the plateau, the special solution is
$\varphi(N)=\varphi_{t}+\frac{\pi_{t}}{3}\left(1-e^{3N}\right),$ (25)
where $\pi_{t}$ is the velocity $\pi\equiv-\mathrm{d}\varphi/\mathrm{d}N$ at the endpoint where $N=0$, such that $\varphi(0)=\varphi_{t}$, $\pi(0)=\pi_{t}$. On the plateau, we have $\pi(N)=\pi_{t}e^{3N}$ as expected, and (25) can be rewritten as a “conservation law”
$\pi(N)+3\varphi(N)=\pi_{t}+3\varphi_{t},$ (26)
which is valid for all the trajectories at any $e$-folding number. For a perturbed universe, (26) gives $\tilde{\pi}(N)+3\tilde{\varphi}(N)=\tilde{\pi}_{t}+3\varphi_{t}$, which in turn gives
$\delta\pi(N)+3\delta\varphi(N)=\delta\pi_{t}.$ (27)
We defined $\delta\varphi=\tilde{\varphi}-\varphi$, $\delta\pi\equiv\tilde{\pi}-\pi$ and $\delta\pi_{t}\equiv\tilde{\pi}_{t}-\pi_{t}$, which are Gaussian variables as $\varphi$ is a massless field. The physical meaning of (26) and (27) can be well understood in the phase portrait Fig. 1.
Figure 1: The schematic phase portrait of the ultra-slow-roll inflation on a flat plateau. The blue and green lines with arrows are the inflaton trajectory in the background universe and a perturbed universe, respectively. The red lines are equal-$N$ lines, determined by the conservation law (26), while ($\tilde{\pi}$, $\tilde{\varphi}$) at the perturbed trajectory should have the same $N$ as ($\pi$, $\varphi$), i.e. the perturbation is along an equal-$N$ line. Such a perturbed point ($\tilde{\pi}$, $\tilde{\varphi}$) satisfies (27), as is clearly shown geometrically. For a detailed discussion of such a phase portrait of ultra-slow-roll inflation, see Passaglia et al. (2019).
From $\pi(N)=\pi_{t}e^{3N}$, we have
$N=\frac{1}{3}\ln\frac{\pi}{\pi_{t}}=-\frac{1}{3}\ln\left(1+\frac{3(\varphi-\varphi_{t})}{\pi}\right),$ (28)
where the second equality comes from (26), which is useful to determine the equal-$N$ lines in Fig.1: They are radial lines from $(\pi=0,~{}\varphi=\varphi_{t})$.
In the separate universe approach, another remote and independent “perturbed universe” has expanded another $e$-folding number of $\tilde{N}=(1/3)\ln\tilde{\pi}/\tilde{\pi}_{t}$, with $\tilde{\pi}=\pi+\delta\pi$, and $\tilde{\pi}_{t}=\pi_{t}+\delta\pi_{t}$. In the $\delta N$ formalism, the late-time curvature perturbation in this perturbed universe is the difference between these $e$-folding numbers
$\delta N=\tilde{N}-N=\frac{1}{3}\ln\frac{\tilde{\pi}}{\tilde{\pi}_{t}}\frac{\pi_{t}}{\pi}=\frac{1}{3}\ln\left(1-\frac{\delta\pi}{\pi}\right)-\frac{1}{3}\ln\left(1-\frac{\delta\pi_{t}}{\pi_{t}}\right)$ (29)
Note that from from (27) we have $\delta\pi\sim\delta\pi_{t}$, so unless $N\approx 0$, the first term in (29) is always neglegible compared to the second term as $\pi=\pi_{t}e^{3N}\gg\pi_{t}$. If the later evolution do not contribute significantly to $\delta N$, [^4] This requirement is not trivial. It is equivalent to require that the inflaton accelerates to the later slow-roll attractor after it classically shoots out of the flat pleateau. On the other hand, if $\varphi$ decelerates, the $\delta N$ contributed from the later stage is dominant, which is usually Gaussian and thus the entire curvature perturbation is nearly Gaussian. See below. the final curvature perturbation is
$\mathcal{R}\approx-\frac{1}{3}\ln\left(1-\frac{\delta\pi_{t}}{\pi_{t}}\right),$ (30)
which is highly non-Gaussian with a nominal nonlinear parameter
$f_{\mathrm{NL}}=\frac{5}{2}.$ (31)
Apparently, the non-Gaussianity expressed in (30) can not be simply described by quadratic expansion with $f_{\mathrm{NL}}=5/2$. We should emphasize that (30) is only valid when the contribution from the later evolution to $\delta N$ is negligible, the meaning of which will be discussed in the subsection below.
III.2 Constant-roll on a quadratic potential
In the previous section we see how the non-attractor solution is important in enhancing the power spectrum and generating large non-Gaussianities. Such a conclusion is general, valid also for other types of potentials. In this subsection we will consider another analytically solvable model – constant-roll inflation on a quadratic potential, which can be inserted as an intermediate stage between two slow-roll stages to enhance the curvature perturbation. We consider the following potential
$V(\varphi)=V_{0}+\frac{1}{2}m^{2}\varphi^{2},$ (32)
where $V_{0}\gg|m^{2}\varphi^{2}|$, such that the Hubble parameter can still be approximated by a constant $H\approx\sqrt{V_{0}}/\sqrt{3}M_{\mathrm{Pl}}$, while the second slow-roll parameter $\eta\equiv m^{2}/3H^{2}$ is also a constant. The equation of motion is
$\frac{{\mathop{}\!\mathrm{d}}^{2}\varphi}{{\mathop{}\!\mathrm{d}}N^{2}}-3\frac{{\mathop{}\!\mathrm{d}}\varphi}{{\mathop{}\!\mathrm{d}}N}+3\eta\varphi=0,$ (33)
which is written in the $e$-folding number $N$, counted backwards in time from a junction point $\varphi_{t}$ far enough from the bump, such that the solution is already an attractor there  [^5] For the curvature perturbation which exits the horizon during the constant-roll stage, there will be still some $e$-folding number from $\varphi_{t}$ to the end of inflation, say $\varphi_{e}$. However, if the trajectories we study here already merged into the attractor solution before $\varphi_{t}$, these Hubble patches evolves uniquely after $\varphi_{t}$, thus the contribution to $\delta N$ from $\varphi_{t}$ to $\varphi_{e}$ is negligible.. Setting $\varphi\propto e^{\lambda N}$, the characteristic root $\lambda$ of (33) is
$\lambda^{2}-3\lambda+3\eta=0,\quad\Longrightarrow\quad\lambda_{\pm}=\frac{3\pm\sqrt{9-12\eta}}{2}.$ (34)
For $\eta<3/4$, we have two real solutions with $\lambda_{-}<\lambda_{+}$, which means $\varphi$ has two independent solutions: the attractor solution $\varphi\propto e^{\lambda_{-}N}$ and the non-attractor solution $\varphi\propto e^{\lambda_{+}N}$. Again, defining the velocity $\pi(N)=-\mathrm{d}\varphi/\mathrm{d}N$, the attractor and non-attractor solutions satisfy
$\pi+\lambda_{\mp}\varphi=0.$ (35)
Considering the boundary condition of $\pi(0)=\pi_{t}$, $\varphi(0)=\varphi_{t}$, the solution is
$\varphi(N)=-\frac{\pi_{t}+\lambda_{-}\varphi_{t}}{\lambda_{+}-\lambda_{-}}e^{\lambda_{+}N}+\frac{\pi_{t}+\lambda_{+}\varphi_{t}}{\lambda_{+}-\lambda_{-}}e^{\lambda_{-}N}.$ (36)
Therefore, we can solve these equations for $N$:
$N=\frac{1}{\lambda_{\pm}}\ln\frac{\pi+\lambda_{\mp}\varphi}{\pi_{t}+\lambda_{\mp}\varphi_{t}}\,,$ (37)
[…]
The two different expressions for $N$ in (37) are equivalent under the “conservation law”
$\left(\frac{\pi+\lambda_{+}\varphi}{\pi_{t}+\lambda_{+}\varphi_{t}}\right)^{\lambda_{+}}=\left(\frac{\pi+\lambda_{-}\varphi}{\pi_{t}+\lambda_{-}\varphi_{t}}\right)^{\lambda_{-}},$ (38)
which is a direct consequence of the solution (36). This is the fiducial $e$-folding number. For a perturbed universe where $\tilde{\varphi}=\varphi+\delta\varphi$, $\tilde{\pi}=\pi+\delta\pi$, $\tilde{\pi}_{t}=\pi_{t}+\delta\pi_{t}$, $\tilde{N}=N+\delta N$, we have
$N+\delta N=\frac{1}{\lambda_{\pm}}\ln\frac{\pi+\delta\pi+\lambda_{\mp}\varphi}{\pi_{t}+\delta\pi_{t}+\lambda_{\mp}\varphi_{t}}\,,$ (39)
Subtracting (37) from (39), we have
$\mathcal{R}=\delta N=\frac{1}{\lambda_{\pm}}\ln\left(1+\frac{\delta\pi+\lambda_{\mp}\delta\varphi}{\pi+\lambda_{\mp}\varphi}\right)-\frac{1}{\lambda_{\pm}}\ln\left(1+\frac{\delta\pi_{t}}{\pi_{t}+\lambda_{\mp}\varphi_{t}}\right)\,.$ (40)
We see that the curvature perturbation has two equivalent expressions, which is called logarithmic duality in Pi and Sasaki (2023b). Apparently it comes from the quadratic nature of the potential, which is good enough when the field excursion we study is not too long. The equivalence of the upper and lower formulas in (40) can be easily shown by the perturbed conservation law (38):
$\left(1+\frac{\delta\pi_{t}}{\pi_{t}+\lambda_{+}\varphi_{t}}\right)^{-\lambda_{+}}\left(1+\frac{\delta\pi_{t}}{\pi_{t}+\lambda_{-}\varphi_{t}}\right)^{\lambda_{-}}=\left(1+\frac{\delta\pi+\lambda_{-}\delta\varphi}{\pi+\lambda_{-}\varphi}\right)^{\lambda_{-}}\left(1+\frac{\delta\pi+\lambda_{+}\delta\varphi}{\pi+\lambda_{+}\varphi}\right)^{-\lambda_{+}}\,.$ (41)
In principle, the perturbation at the boundary $\delta\pi_{t}$ induced by the initial perturbation ($\delta\varphi$, $\delta\pi$) should be (38). However, in this case we do not need it after assuming that the inflaton is already the attractor solution at $\varphi_{t}$, i.e. $\pi_{t}+\lambda_{-}\varphi_{t}\approx 0$ and $\delta\pi_{t}\approx 0$. Therefore, it is more convenient to use the lower line of (40), as the second logarithm is negligible:
$\mathcal{R}\approx\frac{1}{\lambda_{-}}\ln\left(1+\frac{\delta\pi+\lambda_{+}\delta\varphi}{\pi+\lambda_{+}\varphi}\right)=\frac{1}{\lambda_{-}}\ln\left(1-\left.\lambda_{-}\frac{\delta\varphi}{\pi}\right|_{\mathrm{attractor}}\right).$ (42)
In the second step we use the fact that $\delta\varphi$ and $\varphi$ obeys the same equation of motion (33) on superhorizon scales (i.e. $k^{2}\ll H^{2}a^{2}$), when the potential is quadratic  [^6] This is a special case of the well know proportionality $\delta\varphi\propto\dot{\varphi}/H$ on superhorizon scales for any potential in single-field inflation. . Therefore, $(\delta\pi+\lambda_{+}\delta\varphi)/(\pi+\lambda_{+}\varphi)$ is a constant along each trajectory, which can be evaluated at a much later time after horizon exit when $\varphi$ is already in the attractor (but still before the boundary) Atal et al. (2019). The phase diagram is shown in Figure 2.
Slow-roll inflation and ultra-slow-roll inflation discussed in Section III.1 are special cases of (40). Slow roll is the attractor solution when $\lambda_{-}\ll 1$, $\lambda_{+}\approx 3$. Expanding (42), we get (21) with $\mathcal{R}_{g}=-\delta\varphi/\pi=H\delta\varphi/\dot{\varphi}$ and $f_{\mathrm{NL}}=-(5/3)\lambda_{-}$. In ultra-slow-roll inflation, the non-attractor solution dominates. Substituting $\lambda_{-}=0$, $\lambda_{+}=3$ into the upper line of Eq. (40), we can get (29).
III.3 A general analysis on single-field inflation
We have shown how a quadratic piece of the potential (32) can give rise to the logarithmic relation of $\mathcal{R}(\delta\varphi,\delta\pi)$, Eq. (40), where $(\delta\varphi,\delta\pi)$ are evaluated on the initial spatially-flat surface after horizon crossing, which are determined by subhorizon quantum fluctuations. In general, such a potential (32) is embedded into a global potential. The initial condition of $(\varphi,\pi)$ are usually provided by a preceding (slow-roll) stage, while after the bump/plateau, the potential is usually followed by another slow-roll potential to end inflation. Before and after the junction point, the attractor solutions are usually different. This can excite the non-attractor solution transiently, which contribute to the total $\delta N$ as we discussed above. All of such contributions should be taken into account to calculate the final $\delta N$.
We consider a piecewise potential consisting of two parabolas:
$V(\varphi)=\left\{\begin{matrix}&\displaystyle V_{0}+\frac{m^{2}}{2}\left(\varphi^{2}-\varphi_{t}^{2}\right),&\text{for~{}}\varphi\leq\varphi_{t};\\ \\ &\displaystyle V_{0}-\frac{m^{\prime 2}}{2}\left(\varphi_{t}-\varphi_{m}\right)^{2}+\frac{m^{\prime 2}}{2}\left(\varphi-\varphi_{m}\right)^{2},&\text{for~{}}\varphi>\varphi_{t}.\end{matrix}\right.$ (46)
where $\varphi_{t}$ is the junction point of the two quadratic potentials with $V(\varphi_{t})=V_{0}$, and $\varphi_{m}$ is the minimum of $V(\varphi>\varphi_{t})$. For simplicity, we set the origin to make $\varphi=0$ the maximum (minimum) of $V(\varphi<\varphi_{t})$, when $m^{2}<0$ ($m^{\prime 2}>0$). This is an extension of the Starobinsky model with a hinged linear potential Starobinsky (1992) to quadratic level, which can be further extended to a multi-exponential potential Domènech et al. (2023). Inflation ends at $\varphi_{f}$ in the second stage. A schematic figure of this piecewise potential is shown in Fig. 3. $V(\varphi)$ is continuous, but $V^{\prime}(\varphi)$ has a jump at $\varphi_{t}$. This means $\varphi^{\prime}$ is continuous, but a change in the slope means that the slow-roll solution for $\varphi<\varphi_{t}$ is not the slow-roll solution for $\varphi>\varphi_{t}$, thus $\varphi$ will decelerate/accelerate to reach the later slow-roll solution, during which the curvature perturbation will be enhanced/suppressed, and non-Gaussianity could become large.
Figure 3: Schematic pictures of the piecewise quadratic potential $V(\varphi)$ in (46), with $m_{1}^{2}<0$ (left) and $m_{2}^{2}>0$ (right). The potential is continuous at the transition point is $\varphi_{t}$, where $V(\varphi_{t})=V_{0}$, but $V^{\prime}(\varphi_{t})$ has a jump. The origin is chosen at the local maximum (upper panel) or minimum (lower panel) of $V(\varphi<\varphi_{t})$, for $m_{1}^{2}<0$ or $m_{1}^{2}>0$, respectively. At $\varphi_{t}$, the potential is continuous, while its slope may not be. The $e$-folding number is also labeled at $\varphi$, $\varphi_{t}$, and $\varphi_{f}$.
By the method used in Section III.2, we can solve $\delta N$ contributed in all the stages, and add together to get the final curvature perturbation Pi and Sasaki (2023b):
$\mathcal{R}$ $=\delta N_{1}\left(\{\varphi,\pi\}\to\{\varphi_{t},\pi_{t}\}\right)+\delta N_{2}\left(\{\varphi_{t},\pi_{t}\}\to\{\varphi_{e},\pi_{e}\}\right)$ (47)
$\delta N_{1}$ $=\frac{1}{\lambda_{\pm}}\ln\left(1+\frac{\delta\pi+\lambda_{\mp}\delta\varphi}{\pi+\lambda_{\mp}\varphi}\right)-\frac{1}{\lambda_{\pm}}\ln\left(1+\frac{\delta\pi_{t}}{\pi_{t}+\lambda_{\mp}\varphi_{t}}\right)$ (48)
$\delta N_{2}$ $=\frac{1}{\lambda^{\prime}_{\pm}}\ln\left(1+\frac{\delta\pi_{t}}{\pi_{t}+\lambda^{\prime}_{\mp}(\varphi_{t}-\varphi_{m})}\right)-\frac{1}{\lambda^{\prime}_{\pm}}\ln\left(1+\frac{\delta\pi_{e}}{\pi_{e}+\lambda^{\prime}_{\mp}(\varphi_{e}-\varphi_{m})}\right)\,,$ (49)
where $\lambda_{\pm}$ and $\lambda^{\prime}_{\pm}$ are the characteristic roots of each piece, defined by (34). The equivalence of upper and lower expressions of each $\delta N_{i}$ are guaranteed by the perturbed conservation law (41) for each piece. Here, $\delta\varphi$ and $\delta\pi$ are the inflaton perturbation and its momentum (i.e. time derivative with respect to $-N$) on an initial spatially-flat slice around ($\varphi$, $\pi$), which should be solved from the equation of motion for $\delta\varphi$, and evaluated at a moment slightly later than the horizon exit. In slow-roll stage, typically $\delta\varphi\approx H/(2\pi)$ and $\delta\pi\approx 0$ as it decays exponentially after horizon exit, yet it is not so simple for non-attractor stages. For a concrete example of calculating $\delta\varphi$ and $\delta\pi$ in ultra-slow-roll inflation, see Jackson et al. (2023); Domènech et al. (2023). After we know $\delta\varphi$ and $\delta\pi$, the corresponding perturbations on all the later boundaries, i.e. $\delta\pi_{t}$, $\delta\pi_{e}$, etc., can be solved by the perturbed conservation law (41) for each piece, although this equation can only be solved numerically except for a few special values of $\lambda_{-}$. As we have shown before, finally all the different trajectories induced by the initial perturbation ($\delta\pi$, $\delta\varphi$) merge to the unique attractor trajectory. Since then, the $e$-folding number only depend on $\varphi$, and $\delta N$ in this stage only depend on the time difference when $\varphi$ falls on to the attractor, which is usually Gaussian. It is straightforward to extend (47) to include more segments of quadratic potentials, and for every segment of the quadratic piece, two logarithms from the early and late boundaries are present, similar to (49).
Figure 4: The schematic potential (left panel) and the phase portrait (right panel) of the “smooth transition” from ultra-slow-roll to slow-roll, with $\eta^{\prime}=-0.01$, $\varphi_{m}=0$. For blue and orange trajectories we set $\pi=9.15,~{}9.05$ at $\varphi=-3$ (in the unit of $H$), respectively. To show the difference of the trajectories clearly, $\delta\pi$ is exaggerated. Equal-$N$ lines are parallel to $\pi\propto-3\varphi$ with almost equal intervals, thus Gaussianity is well preserved.
As a simple example, let’s discuss ultra-slow-roll inflation on a flat potential connected to a slow-roll potential. Then (48) reduces to (30), while (49) reduces to (42):
$\mathcal{R}\approx-\frac{1}{3}\ln\left(1+\frac{\delta\pi_{t}}{\pi_{t}}\right)+\frac{1}{\lambda^{\prime}_{-}}\ln\left(1+\frac{\delta\pi_{t}}{\pi_{t}+\lambda^{\prime}_{+}\left(\varphi_{t}-\varphi_{m}\right)}\right).$ (50)
In general, the final curvature perturbation is a sum of these two terms, and the PDF of $\mathcal{R}$ is complicated when two logarithmic functions are equally important. However, there are two important limits, which can be analytically solved. When $\pi_{t}\ll\lambda_{-}^{\prime}\left(\varphi_{t}-\varphi_{m}\right)$, the velocity at the end of flat plateau is not yet the attractor of the slow-roll potential, thus a transient acceleration stage is followed. The first term of (50) dominates, so the final curvature perturbation is non-Gaussian with $f_{\mathrm{NL}}=5/2$. On the contrary, when $\pi_{t}\gg\lambda_{-}^{\prime}(\varphi_{t}-\varphi_{m})$, the velocity at the end of flat plateau surpasses the attractor velocity of the slow-roll potential, thus a transient deceleration stage is followed. The second term of (50) dominates, so the final curvature perturbation is nearly Gaussian with $f_{\mathrm{NL}}=-(6/5)\lambda_{-}^{\prime}\approx-(6/5)\eta^{\prime}$. These two limits are called “smooth transition” and “sharp transition” respectively in Cai et al. (2018), where a parameter $h\equiv 6|\lambda^{\prime}_{-}(\varphi_{t}-\varphi_{m})/\pi_{t}|$ was defined to compare the slow-roll attractor velocity $-\lambda^{\prime}_{-}(\varphi_{t}-\varphi_{m})$ and the end-of-ultra-slow-roll velocity $\pi_{t}$ at the transition point $\varphi_{t}$. To avoid confusion, we should emphasize that the comparison is between the velocities, but not between the slope of the potential.
This result is easy to understand by looking at the phase portrait. For the “smooth transition” case, although the potential behaves differently for $\varphi<\varphi_{t}$ and $\varphi>\varphi_{t}$, but the non-attractor solution ($\pi=-3\varphi$) in the flat plateau is almost the same as the non-attractor solution in a slow-roll potential ($\pi=-\lambda_{+}^{\prime}\varphi\approx(3-\eta^{\prime})\varphi$). Therefore, this case is similar to the bumpy potential we studied in Section III.2, and (42) tells us $\mathcal{R}$ is nearly Gaussian with $f_{\mathrm{NL}}=-(6/5)\lambda_{-}^{\prime}$. The phase portrait Fig. 4 looks similar to that of a slow-roll inflation, as the ultra-slow-roll trajectories are similar to the fast-roll trajectories in an ordinary slow-roll inflation, which are parallel to the equal-$N$ lines. The difference in the $e$-folding number for two nearby trajectories are proportional to the difference in $\varphi$, therefore $\delta N\propto\delta\varphi$ which is a Gaussian relation. The well studied Starobinsky model Starobinsky (1992); Leach et al. (2001); Biagetti et al. (2018); Tasinato (2021); Pi and Wang (2023); Domènech et al. (2023) belongs to this limit.
Figure 5: The schematic potential (left panel) and the phase portrait (right panel) of the “sharp transition” from ultra-slow-roll to slow-roll, with $\eta^{\prime}=0.01$, $\varphi_{m}=100$ (in the unit of $H$). For the fiducial (blue) and perturbed (orange) trajectories we set $\pi=9.15,~{}9.05$ at $\varphi=-3$, respectively. To show the difference of the trajectories clearly, $\delta\pi$ is exaggerated. Near the origin, the trajectories merge to a unique fast-roll trajectory with different $\pi$, which contribute negligibly to $\delta N$.
For the sharp transition case, the inflaton experiences an acceleration after the ultra-slow-roll stage, therefore the nearby trajectories, slightly shifted by the velocity difference $\delta\pi_{t}$ generated during the ultra-slow-roll stage, can also induce difference in $e$-folding number $\delta N$ in the later slow-roll stage. However, as $\pi_{t}\ll 1$, trajectories with different $\pi_{t}$’s near the origin are almost on the same fast-roll trajectory in the slow-roll era for $\varphi>\varphi_{t}$, as the trajectories near the $\varphi$-axis are perpendicular to it. On this trajectory, the difference in $N$ caused by $\delta\pi_{t}$ is negligible, and $\varphi=\varphi_{t}$ is approximately an equal-$N$ surface. Following the discussion in Section III.1, we will arrive at the non-Gaussian relation (30) straightforwardly.
III.4 The curvaton scenario
The curvaton scenario is a special two-field inflation model, of which the curvature perturbation is generated by the density perturbation of the curvaton field $\chi$ at its decay after inflation. This was first proposed in Moroi and Takahashi (2001); Enqvist and Sloth (2002); Lyth et al. (2003) as an alternative method of generating curvature perturbation, which was usually referred to as the outcome of the quantum fluctuation of the inflaton field $\varphi$ during inflation. In the simplest curvaton models the power spectrum is nearly scale-invariant. To generate PBHs, the power spectrum should be enhanced on a specific scale, which can be realized in the axion-curvaton model Kasuya and Kawasaki (2009); Kawasaki et al. (2013a, b); Ando et al. (2018b), non-minimal curvaton scenario Pi and Sasaki (2023a); Meng et al. (2023); Chen et al. (2023), or the multi-curvaton model Suyama and Yokoyama (2011). In all the cuvaton models, it was shown that non-Gaussianity is inevitably enhanced compared to the slow-roll inflation, and the non-linear parameter is Bartolo et al. (2004); Enqvist and Nurmi (2005); Sasaki et al. (2006)  [^7] For simplicity, we only consider a quadratic potential for the curvaton $\chi$, and the decay is supposed to be instantaneous. For more details of these assumptions, see Sasaki et al. (2006).
$f_{\mathrm{NL}}=\frac{5}{4r},\qquad r\equiv\frac{3\Omega_{\chi,\mathrm{dec}}}{4-\Omega_{\chi,\mathrm{dec}}},$ (51)
where $r$ is approximately the cosmological parameter of the curvaton $\chi$ at its decay. Apparently such a nominal $f_{\mathrm{NL}}$ is not enough to calculate the PBH abundance, as $f_{\mathrm{NL}}$ is always larger than 1.
We will skip the model building and the dynamics of the curvaton field, but focus on its fully non-Gaussian curvature perturbation on the uniform-density slice, which is derived in Sasaki et al. (2006) for a Gaussian curvaton field perturbation in the sudden decay approximation:
$\zeta=\ln\left(K^{1/2}\frac{1+\sqrt{ArK^{-3/2}-1}}{(3+r)^{1/3}}\right),$ (52)
where $r$ is defined in (51), and
[…]
$P$ $=(Ar)^{2}+\sqrt{(Ar)^{4}+(3+r)(1-r)^{3}},$ (54)
$K$ $=\frac{1}{2}\left[P^{1/3}+(r-1)(3+r)^{1/3}P^{-1/3}\right].$ (55)
This fully non-linear curvature perturbation $\zeta$ can be used to calculate the PBH formation, following the procedure we show in Eq. (20) Ferrante et al. (2022, 2023); Gow et al. (2023).
The fully non-Gaussian curvature perturbation (52) seems complicated, but it has two interesting limiting cases for $r\ll 1$ and $r\to 1$, both of which have fruitful physical implications. In the former limit, the curvaton $\zeta$ is negligible at its decay, we have $r\ll 1$, and (52) can be expanded for small $r$:
$\zeta=\frac{r}{3}(A-1)-\frac{r^{2}}{9}(A-1)^{2}+\cdots=\frac{2}{3}r\frac{\delta\chi}{\chi}+\frac{r}{3}\left(1-\frac{4r}{3}\right)\left(\frac{\delta\chi}{\chi}\right)^{2}+\cdots,$ (56)
where dots are higher order terms of $\mathcal{O}(r^{3})$. We see that $\zeta$ has the familiar form of the quadratic local non-Gaussianity (29), with $\zeta_{g}=(2/3)r\delta\chi/\chi$ and $f_{\text{NL}}=5/(4r)$. Cubic and higher terms are suppressed in the $r\to 0$ limit, which makes it possible to truncate (56) at quadratic level when $f_{\mathrm{NL}}\gg 1$ (i.e. $r\ll 1$). Therefore, counterintuitively, such a quadratic expression (56) holds for large $f_{\mathrm{NL}}$, towards the limit of a $\chi^{2}$-distribution.
The opposite limit $r\to 1$ is more interesting and attracts much attention recently. In this limit, we simply have $P=2(Ar)^{2}$, $K=(Ar)^{2/3}/2$, and
$\zeta=\frac{2}{3}\ln A=\frac{2}{3}\ln\left(1+\frac{\delta\chi}{\chi}\right).$ (57)
A simple derivation of (57) is as follows. As curvaton dominates the universe at its decay, the curvature perturbation on the uniform-curvaton slice $\zeta_{\chi}$ equals to the curvature perturbation on uniform density slice $\zeta$, which gives Pi and Sasaki (2023a)
$\zeta=\zeta_{\chi}(t,\mathbf{x})=\int^{\rho_{\chi}(t,\mathbf{x})}_{\rho_{\chi}(t)}\frac{\mathrm{d}\widetilde{\rho}_{\chi}}{3\widetilde{\rho}_{\chi}}=\frac{1}{3}\ln\frac{\rho_{\chi}(t,\mathbf{x})}{\rho_{\chi}(t)}=\frac{2}{3}\ln\left(1+\frac{\delta\chi}{\chi}\right),$ (58)
where in the last step we used $\rho_{\chi}=\frac{1}{2}m_{\chi}^{2}\chi^{2}$. Therefore, in the $r\to 1$ limit, the curvature perturbation $\zeta$ has a logarithmic dependence on the curvaton perturbation $\delta\chi$, which is similar to (42), but with a negative nonlinear parameter
$f_{\mathrm{NL}}=-\frac{5}{4}.$ (59)
This means that the curvaton scenario can suppress the formation of PBHs in the $r\to 1$ limit, which is a natural model to avoid the PBH overproduction problem when interpreting the nHz stochastic GWs as induced GWs Franciolini et al. (2023a).
IV PBH abundance with non-Gaussian curvature perturbation
The non-Gaussian curvature perturbation discussed in the previous section may have a significantly different tail from the Gaussian tail in the PDF in the large $\mathcal{R}$ regime. As we commented, the quantum fluctuations of the inflaton field on subhorizon scales naturally provide a random variable $\delta\varphi$ which has a Gaussian distribution among all the Hubble patches, as the quantum fluctuation is usually Gaussian, especially when the potential is approximately quadratic. Usually, one can define the Gaussian component of the curvature perturbation $\mathcal{R}_{\mathrm{g}}$. In most of the cases, it can be defined as proportional to the inflaton field perturbation $\delta\varphi$ which recovers the leading order Gaussian relation $\mathcal{R}_{\mathrm{g}}\equiv-(H/\dot{\varphi})\delta\varphi$. As we have discussed in Section II.2, the nonlinear relation $\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})$ can bring non-Gaussianities to the PDF of $\mathcal{R}$, as is shown in (17). Then the PBH abundance should be calculated following (20), by transferring the PDF $\mathbb{P}(\mathcal{R})$ to that of the linear compaction function $\mathbb{P}(\mathscr{C}_{\ell})$.
For instance, in ultra-slow-roll inflation, the nonlinear relation $\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})$ is given by (30):
$\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})=-\frac{1}{3}\ln\left(1-3\mathcal{R}_{\mathrm{g}}\right),\quad$ (60)
where $\mathrm{with}~{}\mathcal{R}_{\mathrm{g}}\equiv\delta\pi_{t}/\pi_{t}$. Then the PDF of $\mathcal{R}$, given by (18), is
$\mathbb{P}(\mathcal{R})=\frac{e^{-3\mathcal{R}}}{\sqrt{2\pi}\sigma_{\mathcal{R}_{\mathrm{g}}}}\exp\left[-\frac{\left(e^{-3\mathcal{R}}-1\right)^{2}}{18\sigma^{2}_{\mathcal{R}_{\mathrm{g}}}}\right].$ (61)
Unlike the Gaussian PDF, it goes as $\exp(-3\mathcal{R})$ when $\mathcal{R}\gg\mathcal{O}(0.1)$, so is called exponential tail PDF. Apprarently, for the bumpy potential (constant-roll) case, we have a similar expression with $\mathbb{P}(\mathcal{R})\propto\exp(\lambda_{-}\mathcal{R})$. Such exponential tails are less suppressed compared to the Gaussian tails, thus the PBH abundance will be greatly enhanced. Although $\mathbb{P}(\mathcal{R})$ is not directly used in the calculation of PBH mass function, as we use $\mathbb{P}(\mathscr{C}_{\ell})$ in the Press-Schechter-type formalism, it is straightforward to see the origin of such enhancement by the tail of $\mathbb{P}(\mathcal{R})$ when $\mathcal{R}\gg\sigma_{\mathcal{R}_{\mathrm{g}}}$.
We will show how to calculate the PBH abundance for non-Gaussian curvature perturbation (60) in the extended Press-Schechter formalism, following Gow et al. (2022). By using (19), The compaction function $\mathscr{C}_{\ell}$ can be written as
$\mathscr{C}_{\ell}\equiv\frac{X}{Y},\qquad\mathrm{with}\quad X\equiv-\frac{4}{3}r\mathcal{R}^{\prime}_{\mathrm{g}}(r),\quad Y\equiv 1-3\mathcal{R}_{\mathrm{g}}.$ (62)
The covariance matrix is $\boldsymbol{\Sigma}=\left(\begin{matrix}\sigma^{2}_{X}&\sigma^{2}_{XY}\\ \sigma^{2}_{YX}&\sigma^{2}_{Y}\end{matrix}\right)$, with
$\sigma^{2}_{X}$ $=\left(-\frac{4}{3}\right)^{2}\int(kr)^{2}\mathcal{P}_{\mathcal{R}\mathrm{g}}(k)\left(\frac{\mathrm{d}j_{0}}{\mathrm{d}z}(kr)\right)^{2}\mathrm{d}\ln k,$ (63)
$\sigma^{2}_{XY}=\sigma^{2}_{YX}$ $=\left(-\frac{4}{3}\right)(-3)\int(kr)\mathcal{P}_{\mathcal{R}\mathrm{g}}(k)j_{0}(kr)\frac{\mathrm{d}j_{0}}{\mathrm{d}z}(kr)\mathrm{d}\ln k,$ (64)
$\sigma^{2}_{Y}$ $=(-3)^{2}\int\mathcal{P}_{\mathcal{R}\mathrm{g}}(k)j_{0}^{2}(kr)\mathrm{d}\ln k,$ (65)
where $j_{0}(x)=\sin z/z$ is the 0th order spherical Bessel function. The joint PDF of $X$ and $Y$ becomes
$\mathbb{P}(X,Y)$ $=\frac{1}{2\pi\sqrt{\mathrm{det}\boldsymbol{\Sigma}}}\exp\left[-\frac{1}{2}(X,1-Y)\boldsymbol{\Sigma}^{-1}\left(\begin{matrix}X\\ 1-Y\end{matrix}\right)\right]$ (68)
$=\frac{1}{2\pi\sigma_{X}\sigma_{Y}\sqrt{1-\varrho^{2}}}\exp\left(-\frac{X^{2}}{2\sigma^{2}_{X}}\right)\exp\left[-\frac{1}{2\left(1-\varrho^{2}\right)}\left(\frac{1-Y}{\sigma_{Y}}-\frac{\varrho}{\sigma_{X}}X\right)^{2}\right],$ (69)
where $\varrho\equiv\sigma_{XY}^{2}/(\sigma_{X}\sigma_{Y})$ is the correlation. Then the PDF of $\mathscr{C}_{\ell}$ reads
$\mathbb{P}(\mathscr{C}_{\ell})=\int\mathrm{d}X\mathrm{d}Y~{}\mathbb{P}(X,Y)\delta\left(\mathscr{C}_{\ell}-\frac{X}{Y}\right)=\int\mathrm{d}X~{}Y\cdot\mathbb{P}\left(X\to\mathscr{C}_{\ell}Y,Y\right).$ (70)
The mass of the PBH obeys a power-law scaling from the critical collapse Choptuik (1993); Evans and Coleman (1994); Koike et al. (1995); Niemeyer and Jedamzik (1998); Hawke and Stewart (2002); Musco et al. (2009)
$\frac{M(\mathscr{C}_{\ell})}{M_{H}}\sim K\left(\mathscr{C}-\mathscr{C}_{\mathrm{th}}\right)^{\gamma}=K\left[\left(\mathscr{C}_{\ell}-\frac{3}{8}\mathscr{C}_{\ell}^{2}\right)-\mathscr{C}_{\mathrm{th}}\right]^{\gamma},$ (71)
where $\gamma\approx 0.36$ and $K\sim 1$. Solving this quadratic algebraic equation of $\mathscr{C}_{\ell}$, we have
$\mathscr{C}_{\ell}=\frac{4}{3}\left(1-\sqrt{1-\frac{3}{2}\mathscr{C}_{\mathrm{th}}-\frac{3}{2}\left(\frac{M}{KM_{H}}\right)^{1/\gamma}}\right).$ (72)
Another solution with $\mathscr{C}_{\ell}>4/3$ is discarded as it is the Type II PBH. With (71), to change the variable from $\mathscr{C}_{\ell}$ to $M$, we have
$\frac{\mathrm{d}\mathscr{C}_{\ell}}{\mathrm{d}\ln M}=\frac{1}{\gamma}\frac{\mathscr{C}_{\ell}-\frac{3}{8}\mathscr{C}_{\ell}^{2}-\mathscr{C}_{\mathrm{th}}}{1-\frac{3}{4}\mathscr{C}_{\ell}}.$ (73)
Therefore, substituting (71), (73) together with (70) back into (20), we get the PBH mass function at the formation
$\beta(M)\mathrm{d}\ln M$ $\approx K\frac{\left(\mathscr{C}_{\ell}-\frac{3}{8}\mathscr{C}_{\ell}^{2}-\mathscr{C}_{\mathrm{th}}\right)^{\gamma+1}}{\gamma\left(1-\frac{3}{4}\mathscr{C}_{\ell}\right)}\mathbb{P}(\mathscr{C}_{\ell})\mathrm{d}\ln M,$ (74)
where $\mathscr{C}_{\ell}$ should be expressed as a function of $M$, given by (72). The PBH mass function today is redshifted from the horizon reentry to equality, as the PBH energy density decays as $a^{-3}$ in the radiation dominated background $\sim a^{-4}$:
$f_{\text{PBH}}(M)$ $\approx 3.81\times 10^{8}\left(\frac{g_{*i}}{106.75}\right)^{-1/4}\left(\frac{h}{0.67}\right)^{-2}\left(\frac{M_{\odot}}{M_{H}}\right)^{1/2}\left(\frac{M}{M_{H}}\right)^{-1/2}\beta(M).$ (75)
The PBH abundance today $f_{\mathrm{PBH}}=\int f_{\mathrm{PBH}}(M)\mathrm{d}\ln M$ over all the mass range, which can be used to determine $\mathcal{P_{R}}$ needed. For PBH as all the dark matter, $f_{\mathrm{PBH}}=1$.
As an example, let’s study the monochromatic power spectrum
$\mathcal{P}_{\mathcal{R}\mathrm{g}}=\mathcal{A_{R}}\delta(\ln k-\ln k_{*}).$ (76)
From (63) to (65), we have
$\sigma_{X}=\frac{4}{3}(k_{*}r)\mathcal{A}_{\mathcal{R}}^{1/2}\left|\frac{\mathrm{d}j_{0}}{\mathrm{d}z}(k_{*}r)\right|,\qquad\sigma_{Y}=3\mathcal{A}_{\mathcal{R}}^{1/2}\left|j_{0}(k_{*}r)\right|,$ (77)
and $\varrho=1$. For simplicity, we choose $r_{m}$ to be the same as the Gaussian case, $k_{*}r_{m}\approx 2.74$. Therefore, the PDF (70) degenerate into one-dimensional, as $X$ and $Y$ are fully correlated by $Y=1-(\sigma_{Y}/\sigma_{X})X$, which, together with $X=\mathscr{C}_{\ell}Y$, gives
$X=\frac{\sigma_{X}\mathscr{C}_{\ell}}{\sigma_{X}+\sigma_{Y}\mathscr{C}_{\ell}},\qquad Y=\frac{\sigma_{X}}{\sigma_{X}+\sigma_{Y}\mathscr{C}_{\ell}},\qquad$ (78)
And the degenerate PDF of $\mathscr{C}_{\ell}$ becomes
$\mathbb{P}(\mathscr{C}_{\ell})=\frac{1}{\sqrt{2\pi\sigma^{2}_{X}}}\left(\frac{\sigma_{X}}{\sigma_{X}+\sigma_{Y}\mathscr{C}_{\ell}}\right)^{2}\exp\left[-\frac{1}{2\sigma^{2}_{X}}\left(\frac{\sigma_{X}\mathscr{C}_{\ell}}{\sigma_{X}+\sigma_{Y}\mathscr{C}_{\ell}}\right)^{2}\right].$ (79)
We calculated the PBH mass function $f_{\mathrm{PBH}}$ for a monochromatic power spectrum in the ultra-slow-roll inflation, shown in the left panel of Fig. 6. The amplitude of the power spectrum of the curvature perturbation $\mathcal{A_{R}}$ is set to $4.151077\times 10^{-3}$, such that $f_{\mathrm{PBH}}=1$, while the $f_{\mathrm{PBH}}(M)$ peaks around $10^{-12}$ solar mass, right in the “asteroid mass window” where PBHs can be all the dark matter. Interestingly, as is clearly seen in Fig. 6, for the narrow peak case, the universal infrared scaling $f_{\mathrm{PBH}}(M)\propto M^{1+1/\gamma}\sim M^{3.78}$ which originates from the critical behavior is sharper than the constraints from the extra-galactic $\gamma$ rays proportional to $M^{3.1\sim 3.4}$ Carr et al. (2010). So the entire asteroid-mass window is open for narrow-peak power spectrum even with non-Gaussianity.
By the same method, for perturbative series of (21), we can calculate the PBH abundance $f_{\mathrm{PBH}}$ as a function of $\mathcal{A_{R}}$ for different $f_{\mathrm{NL}}$’s. This is shown in the right panel of Fig. 6 together with the ultra-slow-roll inflation. We see clearly that a positive $f_{\mathrm{NL}}$ systematically enhance the PBH abundance, which makes the $\mathcal{A_{R}}$ required to realize $f_{\mathrm{PBH}}=1$ smaller, as we commented in the introduction. Also, because of the higher order terms, ultra-slow-roll inflation with fully nonlinear relation (60) requires slightly larger power spectrum compared to the quadratic expansion, although both of them have $f_{\mathrm{NL}}=5/2$. The $\mathcal{A_{R}}$ required to realize PBH dark matter for $M_{k_{*}}=2\times 10^{21}~{}\text{g}$ are listed in Table 1.
Figure 6: Left: The PBH mass function $f_{\mathrm{PBH}}(M)$ given by (75) with a monochromatic power spectrum (76) in a ultra-slow-roll inflation with a sharp end, together with the observational constraints from Carr et al. (2020). The total power and central mass are $\mathcal{A_{R}}=4.151\times 10^{-3}$ and $M_{k_{*}}=2\times 10^{21}~{}\text{g}$, such that $\int f_{\mathrm{PBH}}(M)\mathrm{d}\ln M=1$. Right: The PBH abundance $f_{\mathrm{PBH}}$ as a function of the total power $\mathcal{A_{R}}$, for a monochromatic power spectrum (76). From right to left, the black, purple dotted, blue dashed, cyan, brown dot-dashed, and gray curves are for Gaussian case, $f_{\mathrm{NL}}=1$, $f_{\mathrm{NL}}=5/2$, ultra-slow-roll inflation, $f_{\mathrm{NL}}=10$, $f_{\mathrm{NL}}=10^{2}$, and $f_{\mathrm{NL}}=10^{3}$, respectively.
$f_{\mathrm{NL}}$ $\mathcal{A_{R}}$
0 $6.635506\times 10^{-3}$
1 $4.804923\times 10^{-3}$
5/2 $4.010815\times 10^{-3}$
5/2 (USR) $4.151077\times 10^{-3}$
10 $2.299748\times 10^{-3}$
50 $8.051903\times 10^{-4}$
$10^{2}$ $4.674072\times 10^{-4}$
$10^{3}$ $6.268813\times 10^{-5}$
Table 1: The total power $\mathcal{A_{R}}$ for a monochromatic power spectrum (76) needed to make PBH to be all the dark matter at $M_{k_{*}}=2\times 10^{21}~{}\text{g}$, with different $f_{\mathrm{NL}}$’s. The fourth line is the ultra-slow-roll model with a logarithmic $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ relation (60), although it also has $f_{\mathrm{NL}}=5/2$.
V Induced Gravitational Waves
The curvature perturbation can source gravitational waves via nonlinear coupling terms in the perturbation theory up to second order. In this paper, we only briefly discuss the impact of the non-Gaussian curvature perturbation on the amplitude (via the normalization by the PBH abundance) as well as the shape of the induced GW spectrum. For a general discussion of scalar-induced GWs, see Witkowski et al. (2022); Domènech (2021) as well as Domènech (2024).
For each polarization, the tensor perturbation obeys the following equation of motion Ananda et al. (2007); Osano et al. (2007); Baumann et al. (2007)
$h_{\mathbf{k}}^{\prime\prime}+2\mathcal{H}h_{\mathbf{k}}^{\prime}+k^{2}h_{\mathbf{k}}=\int\frac{d^{3}p}{(2\pi)^{3/2}}e^{ij}_{\mathbf{k}}p_{i}p_{j}\left[2\Phi_{\mathbf{p}}\Phi_{\mathbf{k}-\mathbf{p}}+\frac{4\left(a\Phi_{\mathbf{p}}\right)^{\prime}\left(a\Phi_{\mathbf{k}-\mathbf{p}}\right)^{\prime}}{3(1+w)a^{2}\mathcal{H}^{2}}\right],$ (80)
where $\Phi=\frac{2}{3}\mathcal{R}$ is the curvature perturbation in longitudinal gauge calculated in the radiation dominated era, and $e^{ij}_{\mathbf{k}}$ is one of the polarization tensor. After solving the equation of motion (80), we can calculate the energy spectrum of the GWs generated at the horizon reentry, and evaluate it until today. Suppose the non-Gaussian curvature perturbation given by the local form (17), $\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})$, can be expanded as
$\mathcal{R}=\mathcal{R}_{g}+F_{\mathrm{NL}}\mathcal{R}_{g}^{2}+G_{\mathrm{NL}}\mathcal{R}_{g}^{3}+H_{\mathrm{NL}}\mathcal{R}_{g}^{4}+I_{\mathrm{NL}}\mathcal{R}_{g}^{5}+\cdots$ (81)
where $F_{\mathrm{NL}}=(3/5)f_{\mathrm{NL}}$, $G_{\mathrm{NL}}=(9/25)g_{\mathrm{NL}}$, etc. The GW spectrum today is given by a redshifted GW spectrum from the matter-radiation equality Cai et al. (2019a); Unal (2019)
$\Omega_{\text{GW}}(f,\eta_{0})h^{2}=1.6\times 10^{-5}\left(\frac{g_{*s}(\eta_{k})}{106.75}\right)^{-1/3}\left(\frac{\Omega_{r,0}h^{2}}{4.1\times 10^{-5}}\right)\Omega_{\text{GW,eq}}(f),$ (82)
where
$\Omega_{\text{GW,eq}}(k)$ $=3\int^{\infty}_{0}\mathrm{d}v\int^{1+v}_{|1-v|}\mathrm{d}u\frac{1}{4u^{2}v^{2}}\left[\frac{4v^{2}-(1+v^{2}-u^{2})^{2}}{4uv}\right]^{2}\left(\frac{u^{2}+v^{2}-3}{2uv}\right)^{4}$
$\quad\cdot\left[\left(\ln\left|\frac{3-(u+v)^{2}}{3-(u-v)^{2}}\right|-\frac{4uv}{u^{2}+v^{2}-3}\right)^{2}+\pi^{2}\Theta\left(u+v-\sqrt{3}\right)\right]$
$\quad\cdot\left(\mathcal{P}_{\mathcal{R}}(uk)+F_{\text{NL}}^{2}\int^{\infty}_{0}\mathrm{d}\nu\int^{1+\nu}_{|1-\nu|}\mathrm{d}\mu~{}\frac{\mathcal{P}_{\mathcal{R}}(\mu uk)\mathcal{P}_{\mathcal{R}}(\nu uk)}{\mu^{2}\nu^{2}}+\cdots\right)$
$\quad\cdot\left(\mathcal{P}_{\mathcal{R}}(vk)+F_{\text{NL}}^{2}\int^{\infty}_{0}\mathrm{d}\lambda\int^{1+\lambda}_{|1-\lambda|}\mathrm{d}\rho~{}\frac{\mathcal{P}_{\mathcal{R}}(\lambda vk)\mathcal{P}_{\mathcal{R}}(\lambda vk)}{\lambda^{2}\rho^{2}}+\cdots\right).$ (83)
Here we omit the contribution from the connected part $\langle\mathcal{R}^{4}\rangle_{c}$, which is negligibly small in most of the cases Adshead et al. (2021).
The energy spectrum of the induced GW (83) can only be calculated for perturbative series like (81). Fortunately, because the induced GWs only depends on the power spectrum of the curvature perturbation $\mathcal{P_{R}}$ (at most $\sim 10^{-2}$ for enhanced peak), the higher order terms like $G_{\mathrm{NL}}\mathcal{R}^{3}$, $H_{\mathrm{NL}}\mathcal{R}^{4}$ contribute negligibly amount to the GW spectrum, as long as these coefficients $G_{\mathrm{NL}}$’s are not too large. Loose bounds for them are $G_{\mathrm{NL}}\ll\mathcal{O}(10^{2})F_{\mathrm{NL}}$, $H_{\mathrm{NL}}\ll\mathcal{O}(10^{4})F_{\mathrm{NL}}$, etc., which are easy to be satisfied for most of the models.
There are two important examples. In the curvaton scenario with negligible curvaton at its decay, the perturbative series of (21) is valid even if $f_{\mathrm{NL}}\gg 1$, as the higher order terms ($G_{\mathrm{NL}}\mathcal{R}^{2}$ et al.) are always highly suppressed. In the ultra-slow-roll inflation with a sharp end, $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ is given by (60), of which the nonlinear parameters are
$F_{\mathrm{NL}}=3/2,\quad G_{\mathrm{NL}}=3,\quad H_{\mathrm{NL}}=27/4,\quad\cdots,$ (84)
which surely guarantees that higher order terms are suppressed. It was proved in Abe et al. (2022) that the contribution to the induced GW from the higher order terms of the logarithmic $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ are indeed negligible.
Figure 7: Left: The spectrum of GWs induced by the enhanced narrow-peak power spectrum (76), with PBH as all the dark matter. The amplitude $\mathcal{A_{R}}$ for different parameters follows Table 1. From left to right, the black, purple dotted, blue dashed, cyan thick, brown dot-dashed, magenta dashed, and gray curves are for Gaussian case, $f_{\mathrm{NL}}=1$, $f_{\mathrm{NL}}=5/2$, ultra-slow-roll model, $f_{\mathrm{NL}}=10$, $f_{\mathrm{NL}}=10^{2}$, and $f_{\mathrm{NL}}=10^{3}$ respectively. The designed power-law integrated sensitivity curve of space-borne interferometers LISA (gray thick) Bartolo et al. (2016), Taiji (gray dashed) Wang and Han (2021), TianQin (gray dotted) Liang et al. (2021), and DECIGO (gray thin) Schmitz (2021) are also shown. Right: Two more realistic signals, the spectrum of GWs induced by ultra-slow-roll inflation with a sharp transition to slow-roll (blue), and a smooth transition to slow-roll, i.e. Starobinsky’s linear potential model (red). The right panel was drawn with the help of the SIGWfast Witkowski (2022) package.
According to the current observational constraints, asteroid mass PBHs can serve as all the dark matter, of which the scalar induced GWs are right in the mHz band to be detected by the space-borne interferometers. So as an concrete example, we calculate the induced GW spectra for monochromatic power spectra (76) with different $f_{\mathrm{NL}}$’s, but normalized such that PBHs of $M\sim 2\times 10^{21}~{}\text{g}$ are all the dark matter. The resultant induced GWs peak in the millihertz band, which are shown in Figure 7 together with the designed sensitivity curves of the space-borne interferometers LISA Amaro-Seoane et al. (2013, 2012, 2017); Barausse et al. (2020), Taiji Ruan et al. (2020), TianQin Luo et al. (2016). The induced GW from Gaussian curvature perturbation with a monochromatic power spectrum was firstly derived numerically in Ananda et al. (2007) and analysically in Kohri and Terada (2018), which displays a resonant peak of logarithmic divergence at $(2/\sqrt{3})k_{*}$, and an infrared scaling of $\propto k^{2}$. As we commented before, positive $f_{\mathrm{NL}}$ can enhance the PBH formation, which requires less $\mathcal{A_{R}}$ for PBHs as all the dark matter, thus less GW spectrum. However, when $f_{\mathrm{NL}}\lesssim 10$, the suppression is smaller than one order of magnitude, which makes the shape of the GW spectrum around the peak similar to the Gaussian one and difficult to distinguish by the space-borne inteferometers. In this case, the GWs beyond the cutoff frequency $2k_{*}$ is purely from the non-Gaussian part, which is a smoking gun of primordial non-Gaussianity Cai et al. (2019a); Abe et al. (2022). Although these contributions are tiny compared with the main peak, they could be probed by the deci-hertz GW detector like BBO ($0.1$ to 1 Hz) Crowder and Cornish (2005); Corbin and Cornish (2006) or DECIGO ($10^{-2}$ to 1 Hz) Kawamura et al. (2006, 2011).
As $f_{\mathrm{NL}}$ continues to increase, the main peak contributed by the Gaussian part decreases, yet the non-Gaussian contribution increases and changes the shape of the spectrum significantly. The non-Gaussian source is a convolution of two spectra with different wavenumbers, which gives a power-law function with a sharp UV cutoff. Therefore, the GW spectrum induced by the non-Gaussian part behaved as if it were induced by a power-law spectrum with a UV cutoff, which makes its infrared scaling $\Omega_{\mathrm{GW}}^{\mathrm{(NG)}}\propto k^{3}$, a property for a broad spectrum Cai et al. (2019b). Also, $\Omega_{\mathrm{GW}}^{\mathrm{(NG)}}$ has a mild bump instead of a sharp peak at $\sim\sqrt{3}k_{*}$, of which the amplitude is $\sim F_{\mathrm{NL}}^{4}\mathcal{A}_{\mathcal{R}}^{4}$. It becomes visible by LISA for $f_{\mathrm{NL}}\gtrsim 10$.
When $f_{\mathrm{NL}}$ becomes very large, the Gaussian part is invisible, while the non-Gaussian part dominates with an upper bound of $\Omega_{\mathrm{GW}}\sim 10^{-11}$, which is above the LISA sensitivity curve Cai et al. (2019a). Therefore, the detectability of scalar-induced GWs is more or less robust against the nonlinear parameter $f_{\mathrm{NL}}$, $\lambda_{-}$ (see (34) and (42)), or $r_{\mathrm{curvaton}}$ (see (51) and (52)), which made it an important scientific goal of LISA Barausse et al. (2020); Arun et al. (2022); Auclair et al. (2023); Bagui et al. (2023), Taiji Ren et al. (2023), and TianQin Liang et al. (2021) to probe the PBH dark matter.
We should warn that the $\delta$-function peak we considered above is not physical. For more realistic models, the enhancement and decay of the power spectrum is milder, usually power law. Especially, in the typical model of ultra-slow-roll inflation, $\mathcal{P}_{\mathcal{R}}(k)$ increases as $\propto k^{4}$ before it reaches the maximum Byrnes et al. (2019). There might be modulated oscillations of period $\sim\pi$ if the transition from slow-roll to ultra-slow-roll is abrupt Starobinsky (1992); Biagetti et al. (2018); Cole et al. (2022); Pi and Wang (2023). The UV behavior of the power spectrum depends on the shape of the second slow-roll potential, usually displaying a power law $\mathcal{P_{R}}\propto k^{n_{s}-1}$. If the ultra-slow-roll stage ends sharply, the power spectrum drops off a little bit with new modulated oscillations with a larger period of $\sim\pi e^{N_{\mathrm{USR}}}$ Pi and Wang (2023). According to our discussion in Section III.3, these oscillations marks the sharp end of ultra-slow-roll stage, which are accompanied with the logarithmic relation (30). Based on the analytical expression of $\mathcal{P}_{\mathcal{R}}(k)$ as well as the statistical properties, we can draw the GW spectrum associated with a fixed PBH abundance. Currently, there is no satisfactory method of calculating PBH abundance for such a broad spectrum with non-Gaussianity, so we choose the same parameters in Table 1 to show schematically the induced GW spectrum in the ultra-slow-roll inflation in the right panel of Fig. 7. For comparison, we also draw the induced GWs from Starobinsky’s linear potential model, with represents the smooth end of ultra-slow-roll stage. We can see that similar to the $\delta$-function peak case, non-Gaussianity enhances the UV part of the induced GW spectrum, which makes the primary peak of the GW spectrum obscure. A drop off with modulated oscillations can be recognized as a smoking gun for the non-Gaussianity in the ultra-slow-roll inflation, which are detectable by DECIGO and BBO when PBHs constitute all the dark matter.
When the transition from ultra-slow-roll to the second slow-roll stage is smooth, the curvature perturbation is dominated by the $\delta N$ contributed by the second slow-roll stage, which gives $f_{\mathrm{NL,2}}=(5/12)(n_{s,2}-1)$, where $n_{s,2}-1$ is the UV slope of $\mathcal{P_{R}}(k)$. In such induced GW spectra, there will be no modulated oscillations with period of $\sim\pi e^{N_{\mathrm{USR}}}$, and $f_{\mathrm{NL}}$ is determined by the UV slope of the GW spectrum, as $\Omega_{\mathrm{GW}}(k)\sim k^{2(n_{s,2}-1)}$. This can also be used to probe non-Gaussianity Atal and Domènech (2021).
In summary, it is clearly shown in Figure 6 and Figure 7, the dependence of the induced GW on the non-Gaussianity is mild, once the PBH abundance is fixed. The impact of non-Gaussianity is mainly on the spectral shape, which could be detected in the future.
VI Conclusion and Discussion
In this paper we review the impact of the primordial non-Gaussinity in the curvature perturbation on the PBH formation as well as on the generation of scalar-induced GWs. We discussed some concrete inflation models to realize these interesting non-Gaussian curvature perturbations, and demonstrate how to calculate the PBH abundance by extended Press-Schechter formalism, as well as the leading order induced GWs. The main conclusion is that positive/negative non-Gaussianity will greatly enhance/suppress the PBH formation, while the induced GWs are only mildly enhanced for both cases. These two phenomena can be crosschecked in different mass windows and different frequency bands to probe the small-scale behavior of the primordial curvature perturbation, which are important scientific goals for many current and future experiments.
We focus on the simplest local-type non-Gaussianity, which gives $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ as a local function of Gaussian variable $\mathcal{R}_{\mathrm{g}}$. Various types of non-Gaussianities like equilateral, orthogonal, or folded shapes are also important, not only on the abundance but also on the clustering Shandera et al. (2013); Young et al. (2016); Tada and Yokoyama (2015); Young and Byrnes (2020); Suyama and Yokoyama (2019); Matsubara et al. (2019); Ragavendra et al. (2020); De Luca et al. (2021). Recently, Ref. Matsubara and Sasaki (2022) studied the bispectrum and trispectrum in the PBH formation systematically, and shows that in the narrow power spectrum case, the nonlinear parameter $f_{\mathrm{NL}}$ entering the calculation of PBH formation should be a linear combination of $f_{\mathrm{NL}}$ of local, equilateral, folded, and orthogonal shapes:
$f_{\mathrm{NL}}^{\mathrm{eff}}\equiv f_{\mathrm{NL}}^{\mathrm{loc}}-3f_{\mathrm{NL}}^{\mathrm{eq}}+3f_{\mathrm{NL}}^{\mathrm{fo}}-9f_{\mathrm{NL}}^{\mathrm{ort}},$ (85)
which implies that other shapes of non-Gaussianity should also be taken into account when calculating the PBH abundance. Such non-Gaussianities are important for non-canonical kinetic terms, like for instance DBI inflation Silverstein and Tong (2004); Alishahiha et al. (2004), k-essence Armendariz-Picon et al. (1999); Garriga and Mukhanov (1999), or G-inflation Kobayashi et al. (2010).
Our starting point is the $\delta N$ formalism, which assume that the Hubble patches we study are far enough from each other, such that each patch can be seen as a local Friedmann universe where the field value (and other perturbed quantities) have randomly distributed initial value among the vast number of such patches, and evolves in the later stage independently. This is called seperate universe approach. Such an approach works quite well in the slow-roll regime. However, in the beginning of ultra-slow-roll stage (marked by $\varphi_{\mathrm{i}}$), the gradient terms of both growing mode and decaying mode are rather important, which give rise to the famous $k^{4}$ growth Leach et al. (2001); Byrnes et al. (2019); Cole et al. (2022). Such a $k$-dependence cannot be recovered by seperate universe approach and the classical $\delta N$ formalism based on it Jackson et al. (2023). Fortunately, such a failure is only serious for $\varphi\lesssim\sqrt{\pi}\varphi_{\mathrm{i}}$ Domènech et al. (2023), which makes the seperate universe approach and the $\delta N$ formalsm still valid at the peak of the enhanced power spectrum, which is located at $\varphi_{\mathrm{p}}\approx\pi\varphi_{\mathrm{i}}$ Pi and Wang (2023). Apparently, if we extend the $\delta N$ formalism to include the spatial curvature on the initial slice taken slightly after horizon exit, the leading order correction of the gradient expansion can be taken into account Artigas et al. .
We did not discuss the interesting case of negative non-Gaussianity, which is connected to Type II PBHs. Simple estimate shows that negative non-Gaussianity will significantly suppress the PBH abundance, which can be calculated in a perturbative manner Young and Byrnes (2013). For fully nonlinear curvature perturbation with a logarithmic dependence on the Gaussian variable, the PDF of $\mathcal{R}$ displays a Gumbel-like distribution (also called double exponential suppression):
$\mathbb{P}(\mathcal{R})\sim\exp\left(-c^{2}e^{2\lambda_{-}\mathcal{R}}\right),$ (86)
where the $r\to 1$ limit of the curvaton scenario corresponds to $\lambda_{-}=3/2$. Direct calculation shows that it can also significantly suppress the PBH formation. This suppression was utilized to avoid the overproduction of sub-solar-mass PBHs when interpreting the nHz stochastoc GW background as the scalar-induced GWs Franciolini et al. (2023a); Liu et al. (2023). Recent study by peak theory indicates that the PBH formation for negative $f_{\mathrm{NL}}$ is quite difficult Kitajima et al. (2021), and numerical simulations show that there is a lower bound $f_{\mathrm{NL}}\gtrsim-1.2$ below which the Type I PBHs can not form, based on a monochromatic power spectrum Escrivà et al. (2022a). Around this regime the profiles of the density peaks become complicated, and the Type II PBH are crucial Uehara et al. (2024), which we did not touch here.
In many models, the non-attractor behavior is excited by the sudden change of the slop of the potential, which itself is continuous. This actually means that the change of slope is realized in an energy scale much larger than $H$, i.e. $V^{\prime\prime}\gg H$. Another class of models display a discontinuity in the potential, which causes a sudden loss of the kinetic energy. Large non-Gaussianity can also be generated in such a case, which gives a highly asymmetric PDF of $\mathcal{R}$ Cai et al. (2022a, b); Kawaguchi et al. (2023). If the velocity is not enough, the inflaton may not overshoot the step in some Hubble patches. Recent study shows that such Hubble patches finally collapses to a PBH, as the energy density in the surrounding area become smaller than this patch and start to compress it. When the PBH from such bubbles are dominant, there is a large non-Gaussianity Escrivà et al. (2023).
In principle, for non-quadratic potential, the equation of motion for the inflaton can not be solved analytically, and the analysis based on quadratic potentials in Section III.3 can not apply. However, inspired by the logarithmic relation of $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ and its exponential-taile PDF (61), it is rational to conjecture a more general PDF of $\mathcal{R}$ directly Nakama et al. (2016); Hooshangi et al. (2022, 2023). For instance, a simple guess is
$\mathbb{P}(\mathcal{R})\propto\exp\left(-c\left|\mathcal{R}\right|^{p}\right),$ (87)
where $c$ is a positive constant. The PDF with $p<1$ is called heavy tailed, which is more effective in generating PBHs than the exponential-tail PDF. This extra efficiency is needed to generate enough supermassive PBHs to seed early galaxies, as the power spectrum of $\mathcal{R}$ on that scale is strongly constraints by $\mu$- and $y$-distortions, $\mathcal{P_{R}}\lesssim 10^{-5}$ Nakama et al. (2018); Carr and Silk (2018); Atal et al. (2020b); Liu and Bromm (2022); Gouttenoire et al. (2023). Such a heavy tail PDF can arise from self-interacting curvaton scenario Hooper et al. (2023), while a lighter tail of $p=3/2$ can be generated in the non-perturbative regime of inflation model with a ˙$\dot{\mathcal{R}}^{4}$ interaction Celoria et al. (2021).
Up to now, all the discussion above are based on the classical $\delta N$ formalism, which allows us to divide the field value $\varphi$ in a Hubble patch into “background” $\varphi_{0}$ and “perturbation” $\delta\varphi$. Such a split becomes difficult (if not completely impossible) when the perturbations are large and overwhelm the background value. As we commented, this happens when the inflaton is nearly static on a flat plateau due to the Hubble friction, while quantum diffusion should be taken into account Starobinsky (1986, 1982); Starobinsky and Yokoyama (1994). Although the exponential-tail PDF like (61) appears as the leading order of the PDF in the semi-classical limit, more complicated and bizarre PDF arises in the quantum regime Bullock and Primack (1997); Vennin and Starobinsky (2015); Pattison et al. (2017); Ezquiaga et al. (2020); Vennin (2020); Figueroa et al. (2021); Pattison et al. (2021); Figueroa et al. (2022); Animali and Vennin (2022); Ezquiaga et al. (2023). This is the topic of the next chapter Vennin and Wands (2024).
Acknowledgement
I would like to thank Guillem Domènech, Cristiano Germani, Misao Sasaki, Teruaki Suyama, and Takahiro Tanaka for useful discussions, and to thank Ao Wang for helping me draw Fig 7. I also thank the hospitality of YITP, Kavli IPMU, IBS-CTPU, and KIAS during my visit when this paper was finalized, and thank all the participants of the 2nd International Workshop on Gravitational Waves and the Early Universe for the intriguing discussions. This work is supported in part by the National Key Research and Development Program of China Grant No. 2021YFC2203004, by Project No. 12047503 of the National Natural Science Foundation of China, by the JSPS Grant-in-Aid for Early-Career Scientists No. JP20K14461, by JSPS KAKENHI No. JP24K00624, and by the World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan.